OpenStudy (k.binks):

Solve 2x – 5y = 1 by the elimination method. 3x – 4y = -2 ?

OpenStudy (danjs):

you want to multiply one of the equations or both by a constant value, so that when you add the equations together, a variable will sum to zero

OpenStudy (danjs):

for instance 2x and 3x multiply first equation by 3 will give you 6x multiply second equation by -2 giving -6x added they go to zero and x drops out

OpenStudy (jdoe0001):

ditto \(\large \begin{array}{llcll} 2x-5y=1&{\color{brown}{ \times 4}}\implies &8x\cancel{-20y}=4\\ 3x-4y=-2&{\color{brown}{ \times -5}}\implies &-15x\cancel{+20y}=10 \\\hline\\ &&\square ?+\square ?=\square ? \end{array}\)

OpenStudy (k.binks):

-7x= 14 x=-2?

OpenStudy (jdoe0001):

yeap

OpenStudy (k.binks):

Thank you both!

OpenStudy (jdoe0001):

yw